OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, –3) with a slope of –3? 3x + y = 12 3x + y = –12 3x – y = 12 3x – y = –12

OpenStudy (anonymous):

Equation is given by: \[\huge y - y_1 = m(x - x_1)\]

OpenStudy (anonymous):

m = -3 and x1 = 5 and y1 = -3, \[y + 3 = -3(x - 5)\] \[y = -3x + 12\] \[3x + y = 12\]